# Explainability of AI Uncertainty: Application to Multiple Sclerosis Lesion Segmentation on MRI

### v1

@@ Abstract
Addresses clinical explainability gaps in medical AI uncertainty methods.
Novel framework explains AI uncertainty in MS lesion segmentation.
Lesion–wise deep ensemble uncertainty is explained through lesion and patient factors.
Explanations correspond to the factors impending expert rater confidence.
Extensive validation across two large cortical lesion MRI datasets.
@@ 1 Introduction
Trustworthy artificial intelligence (AI) refers to AI systems designed and implemented to meet essential ethical and performance standards, ensuring safe and beneficial integration into society [1, 2]. Ensuring AI trustworthiness is crucial in high-risk fields such as healthcare, where AI-driven decisions directly impact patient outcomes [3, 4]. Regulatory efforts, including the European Union’s AI Act [5], emphasize transparency, accountability, and reliability, mandating clear documentation and interpretability of AI systems to safeguard patient welfare and rights. Recent international guidelines for AI in healthcare, FUTURE-AI [6], have highlighted fairness, universality, traceability, usability, robustness, and explainability as key characteristics needed in the development and deployment of trustworthy AI in healthcare.
@@
Explainable AI (XAI) methods [7] and uncertainty quantification (UQ) [8] play a critical role in trustworthy AI by directly addressing several key attributes outlined by international guidelines and regulatory bodies. XAI methods primarily improve traceability, usability, and explainability by providing insights into how models arrive at specific predictions [7, 6]. UQ significantly contributes to robustness, usability, and explainability by quantifying the reliability of model predictions, highlighting areas of potential model failure, and assisting users in interpreting ambiguous outcomes [8, 6]. Together, UQ and XAI methods facilitate informed interactions between humans and AI, promoting trust and supporting robust clinical decision-making [9].
@@
In this context, UQ has become increasingly important in medical image analysis, particularly in tasks involving deep learning (DL) based segmentation of pathological features [10, 4, 11]. UQ assists clinicians by identifying ambiguous or unreliable predictions, guiding them to areas needing manual review, thus improving clinical confidence and decision quality. However, effectively evaluating uncertainty is inherently challenging due to the absence of definitive ground truth and the diversity of methods used to quantify uncertainty [8]. Although substantial technical progress has been achieved, the clinical impact of UQ remains underinvestigated [12, 13]. Common evaluation strategies assess the relationship of quantified uncertainty with error but do not reveal whether uncertainty is associated with clinically meaningful factors (e.g., lesion characteristics, image quality) or if it aligns with the expert definition of confidence [4]. To bridge this gap, UQ needs to be evaluated not only for technical performance but also for clinical informativeness and capacity to improve human-AI collaboration [14, 9].
@@
In this work, we explore the trustworthy behaviour of DL-based neuroimage analysis to support the assessment of multiple sclerosis (MS) [15, 16]. These patients suffer from a chronic neurodegenerative disease, subsequently evaluated with the use of magnetic resonance imaging (MRI) biomarkers that can be automatically assessed with ad-hoc AI systems. The availability of recommendations and guidelines makes the automated segmentation of MS lesions an ideal scenario to further explore the expert explainability [16, 15, 17]. Focal lesion segmentation in MS serves as an example application to study UQ and explainability, given its inherent complexity and multiple sources of uncertainty [18]. Model uncertainty arises from limited training data and distributional shifts, whereas the irreducible data uncertainty is driven by the data noise and the ambiguity of the definition of the pathology [8]. This complexity intensifies for recently incorporated biomarkers in the diagnostic criteria for MS, such as cortical lesions (CLs), paramagnetic rim lesions, and central vein sign, where sources of uncertainty are magnified due to their small size, subtle contrast differences, and complexity in segmentation [19]. Thus, AI models that can reliably quantify these advanced biomarkers and integrate them safely into the patient assessment workflow are needed to diagnose, treat, and understand the disease course.
@@ 1.1 Contributions
In this study, we propose a novel framework to explain AI uncertainty beyond its relationship with prediction errors, specifically within the context of CL segmentation in MS neuroimaging. The proposed framework is used to evaluate the statistical dependencies between deep ensemble lesion-wise uncertainty and both lesion-specific and patient-specific features. Our framework aims to: (i) explain model behavior, (ii) perform uncertainty sanity checks, and (iii) differentiate between informative uncertainty that supports decision-making and misleading uncertainty that could adversely influence clinical interpretations. Through expert rater feedback, we investigated factors influencing annotator confidence and explored their intersection with high-uncertainty factors. We conducted the analysis across two comprehensive datasets, containing 206 patients and approximately 2000 CLs, evaluated both in-domain and under distribution shifts, extending our previous work in an important way [20].
@@ 2.1 Lesion segmentation in MS
MS is a chronic inflammatory and neurodegenerative disease with a prevalence of 2 million worldwide, characterized by focal, inflammatory demyelination in the central nervous system [15, 16]. MRI plays a key role in MS diagnosis, prognosis, and treatment planning [17]. White matter lesions (WML) detectable on MRI are a hallmark of MS, included in the diagnostic criteria and follow-up pipelines [16]. DL models have been widely explored to automate the tedious and time-consuming annotation process of WML annotation on MRI [21, 22, 23]. Several studies  [23, 24] have highlighted the utility of U-Net-like architectures  [25, 26], dominated by nnU-Net [27, 28] in recent years. CLs have emerged as a promising MRI biomarker that supports the differential diagnosis of MS and aids in predicting disease progression [16, 29, 19, 30, 31]. The CL segmentation on MRI presents a greater challenge both for radiologists and machine learning models, given their small size (as small as a few voxels or milliliters), ambiguity in CL definition and confusion with the WML juxtacortical lesions (see Figure 7). Due to the lack of publicly available data and the complexity of the task, the automation of the CL segmentation task has not been widely explored. Some articles have proposed joint segmentation of CLs and WMLs [32, 33], for which the performance can be dominated by the more prevalent and visible WMLs. CL segmentation was explored in the context of ultra-high field MRI [34] and comparing the visibility of CL on several advanced MRI modalities [35].
@@ 2.2 Uncertainty quantification for medical image segmentation
UQ is widely adopted within the general field of medical image segmentation [10, 4, 11]. A range of methodological approaches have been proposed, including Bayesian methods, ensemble-based methods, conformal prediction and others (such as test-time augmentations, heteroscedastic models, and uncertainty prediction) [10, 4, 11]. Methods such as Monte Carlo dropout [36] and deep ensembles [37] dominate the landscape, with ensembles providing superior uncertainty calibration, albeit at a higher computational cost [4]. To mitigate this, single-inference approaches were explored, including temperature-scaled softmax probability calibration, evidential DL-based on Dempster–Shafer theory, and learned uncertainty frameworks, predicting uncertainty as one of the outputs [4]. In parallel, generative models—such as probabilistic U-Nets [38], PHISeg [39], or diffusion-based architectures [40] — were employed to sample diverse plausible segmentation masks, enabling fine-grained modeling of aleatoric uncertainty and inter-rater variability, although often at a higher computational cost. Test-time augmentation was proposed within the medical imaging field as a model-agnostic way to quantify aleatoric uncertainty [41]. Several approaches go beyond per-voxel uncertainty and propose to quantify uncertainty per region of interest (ROI), e.g., tumor [42] or whole brain [43, 44]. Commonly, these uncertainty measures average the voxel-wise uncertainty within a ROI [45, 42, 43] or quantify the discrepancies in the binary ROI prediction across different samples [42, 44].
For the task of WML MS segmentation trustworthiness assessment using UQ was previously explored using different techniques, including label-flip uncertainty prediction method [46, 47], Monte Carlo dropout [45, 48], batch ensembles [47] and deep ensembles [48, 44]. Several studies proposed to quantify uncertainty associated with each predicted lesion region to aid in the false positive lesion discovery[45, 49] or segmentation quality prediction[50]. Further usage of uncertainty varies, including trustworthiness assessment, quality control, data correction, active learning, and out-of-distribution detection [10, 4, 11].
@@ 2.3 Evaluating uncertainty quantification
UQ evaluation is an active area of research [51]. Assessing the relationship between uncertainty and error is a common practice, however, it does not unravel all the information brought by the uncertainty values [52]. A common evaluation pipeline relies on the assessment of uncertainty calibration (e.g., expected calibration error) [53] or correlation with error [42], robustness-uncertainty measures (e.g., error retention curves) [54, 55, 44]. Some works compare uncertainty against the inter-rater disagreement for segmentation tasks [38, 39]. Qualitative visual assessment of the uncertainty maps serves as a sanity check and enhances the interpretability of the uncertainty values. For instance, high uncertainty values are known to appear at the borders of ROIs [39, 46, 55, 44].
Clinical evaluation of uncertainty is limited to the perceptual assessment rather than clinical informativeness analysis [12, 13]. A pioneering study  [12] on the XAI and uncertainty perception in the clinics found that trust scores were the most positively perceived explanation type among pathologists, as they provided intuitive and immediately actionable confidence information. A more recent work  [13] studies the visualization scenarios of uncertainty maps in the clinics. This work reported that clinicians preferred binary structure-wise interpretable uncertainty visualizations, paired with voxel-level detail when needed.
Several studies have proposed explainability approaches to unravel the information behind the uncertainty values. One study on diabetic retinopathy images classification [56] analyzed the causes of uncertainty by examining the predictive variance across Monte Carlo dropout samples and visualizing its behavior to clinical evaluation or unfamiliar inputs, offering qualitative explanations of model uncertainty. Another study [57] used t-SNE to explain uncertainty in Bayesian CNN-based classification of breast histopathology images, showing how high-uncertainty inputs clustered in latent space and could be flagged for human review, while low-uncertainty cases exhibited significantly improved classification accuracy. To the best of our knowledge, the explainability of uncertainty has not been explored for medical image segmentation.
@@ 3.1 Data
We used two datasets acquired at the i) University Hospital Basel [58] and ii) Lausanne University Hospital [59] in Switzerland. In total, we included data from 206 patients diagnosed with MS according to McDonald criteria [16], having different clinical phenotypes: relapsing-remitting (140), primary-progressive (22), and secondary-progressive (44). The dataset comprises 3D magnetization-prepared 2 rapid gradient-echo (MP2RAGE) images [60] and manual CL annotations. In University Hospital Basel, the imaging was performed using a standardized acquisition protocol on a 3 Tesla whole-body MR system (MAGNETOM Prisma, Siemens Healthcare) with a 64-channel head and neck coil. In Lausanne University Hospital, a 3 Tesla whole-body MR system was used (MAGNETOM Trio, Siemens Healthcare) with a 32-channel head and neck coil.
In both medical centers, annotations were obtained by the consensus: in University Hospital Basel, by a neurologist and a medical doctor with 5 and 5 years of experience in MRI; in Lausanne University Hospital, by a neurologist and a radiologist with 11 and 7 years of experience in MRI. The identified CLs encompassed both intracortical lesions (confined within the gray matter (GM)) and leukocortical lesions (involving both the cortical GM and the adjacent white matter (WM)).
The data from the University Hospital (n=163) was split into training, validation, and test sets with 109:13:41 subjects respectively, corresponding to 857:69:301 CLs. The data from Lausanne University Hospital was solely used for the out-of-distribution evaluation and contains 758 CLs. Further in the manuscript, the in-domain and out-of-domain test sets will be referred to as Test-in and Test-out. The distibutional shift between Test-in and Test-out is due to the change of the medical center, MRI scanner model, acquisition protocol, annotators, annotation protocols, and lesion intensity distributions. Examples of CLs are shown in Figure 1. The summary of the data information is given in Table 1.

### v2

# Explaining Uncertainty in Multiple Sclerosis Lesion Segmentation Beyond Prediction Errors

@@ Abstract
Framework explains lesion-level uncertainty using clinical and imaging features.
Uncertainty is linked not only to error but also to meaningful clinical factors.
Domain shift affects uncertainty behavior and interpretability.
Expert feedback helps assess the clinical relevance of uncertainty explanations.
The method is generalizable across tasks, modalities, and clinical settings.
@@ 1 Introduction
Trustworthy artificial intelligence (AI) refers to AI systems designed and implemented to meet essential ethical and performance standards, ensuring safe and beneficial integration into society  (Thiebes et al., 2021; Li et al., 2023). Ensuring AI trustworthiness is crucial in high-risk fields such as healthcare, where AI-driven decisions directly impact patient outcomes  (Albahri et al., 2023; Lambert et al., 2024). Recent international guidelines for AI in healthcare, FUTURE-AI (Lekadir et al., 2025), have highlighted fairness, universality, traceability, usability, robustness, and explainability as key characteristics needed in the development and deployment of trustworthy AI in healthcare. Regulatory efforts, including the European Union’s AI Act (Union, 2025), emphasize transparency, accountability, and risk management, mandating clear documentation and interpretability of AI systems to safeguard patient welfare and rights.
@@
Explainable AI (XAI) methods  (Minh et al., 2022) and uncertainty quantification (UQ)  (Gawlikowski et al., 2023) play a critical role in trustworthy AI by directly addressing several key attributes outlined by international guidelines and regulatory bodies. On the one side, XAI methods primarily improve traceability, usability, and explainability by providing insights into how models arrive at specific predictions  (Minh et al., 2022; Lekadir et al., 2025). With this, XAI methods aid AI systems to comply with the transparency requirement (Art. 13) of the European Union’s AI Act (Union, 2025). On the other side, UQ significantly contributes to robustness, usability, and explainability by quantifying the reliability of model predictions, highlighting areas of potential model failure, and assisting users in interpreting ambiguous outcomes  (Gawlikowski et al., 2023; Lekadir et al., 2025). These properties show the potential of UQ to adhere to the risk management (Art. 9), human oversight (Art. 14), and robustness (Art. 15) legal requirements (Union, 2025). Together, UQ and XAI methods facilitate informed interactions between humans and AI, promoting trust and supporting robust clinical decision-making  (Zhang et al., 2020; Salvi et al., 2025). However, when applied separately, XAI and UQ do not fully address the comprehensive requirements of Trustworthy AI (Union, 2025; Lekadir et al., 2025).
@@
During the past years, UQ has become increasingly important in medical image analysis, particularly in tasks involving deep learning (DL) based segmentation of pathological features  (Zou et al., 2023; Lambert et al., 2024; Huang et al., 2024). UQ assists clinicians by identifying ambiguous or unreliable predictions, guiding them to areas needing manual review, thus improving clinical confidence and decision quality. However, effectively evaluating uncertainty is inherently challenging due to the absence of definitive ground truth and the diversity of methods used to quantify uncertainty  (Gawlikowski et al., 2023). Although substantial technical progress has been achieved, the clinical impact of UQ remains underinvestigated  (Evans et al., 2022; Huet-Dastarac et al., 2024). Common evaluation strategies assess the relationship of quantified uncertainty with error but do not reveal whether uncertainty is associated with clinically meaningful factors (e.g., lesion characteristics, image quality) or if it aligns with the expert definition of confidence  (Lambert et al., 2024). Furthermore, uncertainty is a non-intuitive concept for humans and does not necessarily correlate linearly with annotators’ confidence (Huet-Dastarac et al., 2024). While the opaque nature of uncertainty hinders clinical adoption, methods that explain the information carried by uncertainty can bridge this gap. Hence, UQ must be evaluated not only for technical performance but also for clinical informativeness and, ideally, for its capacity to improve human-AI collaboration (Lee and See, 2004; Zhang et al., 2020).
@@
This work proposes an interpretability layer integrated into a standard UQ pipeline, enhancing the clinical informativeness of uncertainty measures. Through this approach, we explore the trustworthy behavior of DL-based neuroimage analysis to support multiple sclerosis (MS) assessment (Reich et al., 2018; Thompson et al., 2018). MS patients suffer from a chronic neurodegenerative disease, subsequently evaluated with the use of magnetic resonance imaging (MRI) biomarkers that can be automatically assessed by dedicated AI systems (Belwal and Singh, 2025; Werthen-Brabants et al., 2025). The availability of recommendations and guidelines makes the automated segmentation of MS lesions an ideal scenario to explore further the expert explainability  (Thompson et al., 2018; Reich et al., 2018; Hemond and Bakshi, 2018). Focal lesion segmentation in MS serves as an example application to study UQ and explainability, given its inherent complexity and multiple sources of uncertainty  (Werthen-Brabants et al., 2025). Model uncertainty arises from limited training data and distributional shifts, whereas the irreducible data uncertainty is driven by the data noise and the ambiguity of the definition of the pathology (H ullermeier and Waegeman, 2021). This complexity intensifies for recently incorporated biomarkers in the diagnostic criteria for MS, such as cortical lesions (CLs), paramagnetic rim lesions, and central vein sign, where sources of uncertainty are magnified due to their small size, subtle contrast differences, and complexity in segmentation  (La Rosa et al., 2022). Therefore, reliable AI models capable of quantifying these advanced biomarkers and safely integrating them into clinical workflows are essential for effective disease diagnosis, treatment, and management (Belwal and Singh, 2025; Werthen-Brabants et al., 2025).
@@ 1.1 Contributions
This study significantly extends our preliminary research (Molchanova et al., 2025), proposing a novel framework for explaining uncertainty in CL segmentation using MRI for MS patient care. Our framework shifts focus beyond simply relating uncertainty to prediction errors, aiming instead to improve medical expert understanding of the uncertainty values through detailed, clinically relevant explanations. Specifically, the proposed framework:
  1. 1.
    Explains the variability in lesion-scale uncertainty in terms of lesion and patient characteristics to unravel the information carried by uncertainty values.
  2. 2.
    Performs sanity checks on uncertainty estimates to confirm their reliability and clinical usefulness in different scenarios: in-domain and under distributional shift.
  3. 3.
    Differentiates between useful uncertainty, which aids clinical decisions, and misleading uncertainty, which might negatively affect clinical judgments.
Compared to our earlier study, we introduced several substantial improvements:
  • 1.
    Clinically meaningful features: New lesion-specific and patient-related features were added to improve clinical relevance and interpretability of uncertainty.
  • 2.
    Dataset expansion: The dataset now includes 206 patients and around 2000 lesions from two distinct clinical centers, significantly improving the robustness and generalizability of the analysis.
  • 3.
    Domain-shift analysis: We expanded the evaluation from a single dataset to multiple centers, demonstrating how uncertainty information changes when tested on data from different medical domains.
  • 4.
    Transferability of explanations: We assessed how well uncertainty explanations derived from one dataset apply to others, enhancing generalizability and clinical applicability.
  • 5.
    Ablation study: Multiple regression methods (ordinary least squares, ElasticNet, random forests) and feature combinations were systematically tested to identify the best approach for explaining uncertainty variability.
  • 6.
    Medical expert validation: We conducted structured feedback sessions with medical experts to identify and confirm factors that affect clinician confidence, linking these insights directly to the factors influencing predictive uncertainty.
@@ 2.1 Lesion segmentation in MS
MS is a chronic inflammatory and neurodegenerative disease with a prevalence of 2 million worldwide, characterized by focal, inflammatory demyelination in the central nervous system  (Reich et al., 2018; Thompson et al., 2018). MRI plays a key role in MS diagnosis, prognosis, and treatment planning  (Hemond and Bakshi, 2018). White matter lesions (WML) detectable on MRI are a hallmark of MS, included in the diagnostic criteria and follow-up pipelines  (Thompson et al., 2018). DL models have been widely explored to automate the tedious and time-consuming annotation process of WML annotation on MRI  (Commowick et al., 2018; Kaur et al., 2021; Spagnolo et al., 2023). Several studies  (Spagnolo et al., 2023; Rondinella et al., 2025) have highlighted the utility of U-Net-like architectures  (Ronneberger et al., 2015; Çiçek et al., 2016), dominated by nnU-Net (Isensee et al., 2021; González et al., 2023) in recent years. CLs have emerged as a promising MRI biomarker that supports the differential diagnosis of MS and aids in predicting disease progression  (Thompson et al., 2018; Madsen et al., 2021; La Rosa et al., 2022; Cagol et al., 2024; Beck et al., 2024). The CL segmentation on MRI presents a greater challenge both for radiologists and machine learning models, given their small size (as small as a few voxels or milliliters), ambiguity in CL definition, and confusion with the WML juxtacortical lesions (see Figure 7). Due to the lack of publicly available data and the complexity of the task, the automation of the CL segmentation task has not been widely explored. Some articles have proposed joint segmentation of CLs and WMLs  (Fartaria et al., 2019; Rosa et al., 2020), for which the performance can be dominated by the more prevalent and visible WMLs. CL segmentation was explored in the context of ultra-high field MRI  (La Rosa et al., 2022) and compared the visibility of CL on several advanced MRI modalities  (Gordaliza et al., 2025).
@@ 2.2 Uncertainty quantification for medical image segmentation
UQ is widely adopted within the general field of medical image segmentation  (Zou et al., 2023; Lambert et al., 2024; Huang et al., 2024). A range of methodological approaches have been proposed, including Bayesian methods, ensemble-based methods, conformal prediction, and others (such as test-time augmentations, heteroscedastic models, and uncertainty prediction)  (Zou et al., 2023; Lambert et al., 2024; Huang et al., 2024). Methods such as Monte Carlo dropout (Gal and Ghahramani, 2016) and deep ensembles (Lakshminarayanan et al., 2017) dominate the landscape, with ensembles providing superior uncertainty calibration, albeit at a higher computational cost (Lambert et al., 2024). To mitigate this, single-inference approaches were explored, including temperature-scaled softmax probability calibration, evidential DL-based on Dempster–Shafer theory, and learned uncertainty frameworks, predicting uncertainty as one of the outputs  (Lambert et al., 2024). In parallel, generative models—such as probabilistic U-Nets (Kohl et al., 2018), PHISeg (Baumgartner et al., 2019), or diffusion-based architectures (Amit et al., 2023) — were employed to sample diverse plausible segmentation masks, enabling fine-grained modeling of aleatoric uncertainty and inter-rater variability, although often at a higher computational cost. Test-time augmentation was proposed within the medical imaging field as a model-agnostic way to quantify aleatoric uncertainty (Wang et al., 2019). Several approaches go beyond per-voxel uncertainty and propose to quantify uncertainty per region of interest (ROI), e.g., tumor  (Roy et al., 2019) or whole brain  (Jungo et al., 2020; Molchanova et al., 2025). Commonly, these uncertainty measures average the voxel-wise uncertainty within an ROI (Nair et al., 2018; Roy et al., 2019; Jungo et al., 2020) or quantify the discrepancies in the binary ROI prediction across different samples (Roy et al., 2019; Molchanova et al., 2025).
For the task of WML MS segmentation, trustworthiness assessment using UQ was previously explored using different techniques, including label-flip uncertainty prediction method  (McKinley et al., 2020; Lambert et al., 2022), Monte Carlo dropout  (Nair et al., 2018; Malinin et al., 022b), batch ensembles  (Lambert et al., 2022), and deep ensembles  (Malinin et al., 022b; Molchanova et al., 2025). Several studies proposed to quantify uncertainty associated with each predicted lesion region to aid in the false positive lesion discovery (Nair et al., 2018; Molchanova et al., 2023) or segmentation quality prediction (Lambert et al., 2022). Further usage of uncertainty varies, including trustworthiness assessment, quality control, data correction, active learning, and out-of-distribution detection  (Zou et al., 2023; Lambert et al., 2024; Huang et al., 2024).
@@ 2.3 Evaluating uncertainty quantification
UQ evaluation is an active area of research  (Pignet et al., 2024). Assessing the relationship between uncertainty and error is a common practice, however, it does not unravel all the information brought by the uncertainty values  (Mukhoti and Gal, 2019). A common evaluation pipeline relies on the assessment of uncertainty calibration (e.g., expected calibration error)  (Guo et al., 2017) or correlation with error  (Roy et al., 2019), robustness-uncertainty measures (e.g., error retention curves)  (Malinin et al., 022a; Mehta et al., 2022; Molchanova et al., 2025). Some works compare uncertainty against the inter-rater disagreement for segmentation tasks  (Kohl et al., 2018; Baumgartner et al., 2019). Qualitative visual assessment of the uncertainty maps serves as a sanity check and enhances the interpretability of the uncertainty values. For instance, high uncertainty values are known to appear at the borders of ROIs  (Baumgartner et al., 2019; McKinley et al., 2020; Mehta et al., 2022; Molchanova et al., 2025).
Clinical evaluation of uncertainty is limited to the perceptual assessment rather than clinical informativeness analysis  (Evans et al., 2022; Huet-Dastarac et al., 2024). A pioneering study  (Evans et al., 2022) on the XAI and uncertainty perception in the clinics found that trust scores were the most positively perceived explanation type among pathologists, as they provided intuitive and immediately actionable confidence information. A more recent work  (Huet-Dastarac et al., 2024) studies the visualization scenarios of uncertainty maps in the clinics. This work reported that clinicians preferred binary structure-wise interpretable uncertainty visualizations, paired with voxel-level detail when needed.
Several studies have proposed explainability approaches to unravel the information behind the uncertainty values. One study on diabetic retinopathy images classification (Leibig et al., 2017) analyzed the causes of uncertainty by examining the predictive variance across Monte Carlo dropout samples and visualizing its behavior to clinical evaluation or unfamiliar inputs, offering qualitative explanations of model uncertainty. Another study (Thiagarajan et al., 2022) used t-SNE to explain uncertainty in Bayesian CNN-based classification of breast histopathology images, showing how high-uncertainty inputs clustered in latent space and could be flagged for human review, while low-uncertainty cases exhibited significantly improved classification accuracy. To the best of our knowledge, the explainability of uncertainty has not been explored for medical image segmentation.
@@ 3.1 Data
We used two datasets: i) from a clinical trial acquired at the University Hospital Basel (Granziera, 2018) and ii) a research dataset acquired at the Lausanne University Hospital (Bonnier et al., 2014) in Switzerland. In total, we included data from 206 patients diagnosed with MS according to McDonald criteria (Thompson et al., 2018), having different clinical phenotypes: relapsing-remitting (140), primary-progressive (22), and secondary-progressive (44). The dataset comprises 3D magnetization-prepared 2 rapid gradient-echo (MP2RAGE) images (Kober et al., 2012) and manual CL annotations. In the University Hospital Basel, the imaging was performed using a standardized acquisition protocol on a 3 Tesla whole-body MR system (MAGNETOM Prisma, Siemens Healthcare) with a 64-channel head and neck coil. In Lausanne University Hospital, a 3 Tesla whole-body MR system was used (MAGNETOM Trio, Siemens Healthcare) with a 32-channel head and neck coil.
In both medical centers, annotations were obtained by consensus: in University Hospital Basel, by a neurologist and a medical doctor with 5 and 5 years of experience in MRI; in Lausanne University Hospital, by a neurologist and a radiologist with 11 and 7 years of experience in MRI. The identified CLs encompassed both intracortical lesions (confined within the gray matter (GM)) and leukocortical lesions (involving both the cortical GM and the adjacent white matter (WM)).
The data from the University Hospital (n=163) was split into training, validation, and test sets with 109:13:41 subjects respectively, corresponding to 857:69:301 CLs. The data from Lausanne University Hospital was solely used for the out-of-distribution evaluation and contains 758 CLs. Further in the manuscript, the in-domain and out-of-domain test sets will be referred to as Test-in and Test-out. The distributional shift between Test-in and Test-out is due to the change of the medical center, MRI scanner model, acquisition protocol, annotators, annotation protocols, and lesion intensity distributions. Examples of CLs are shown in Figure 1. The summary of the data information is given in Table 1.